OpenStudy (anonymous):

Simplify 53√96 Answers: 2^3√12 20^3√6 7^3√12 10^3√12

OpenStudy (anonymous):

is this \[5\sqrt[3]{96}\]

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

22.89

OpenStudy (anonymous):

\[96=2^5\times 3\]

OpenStudy (anonymous):

so you get \[5\times 2\sqrt[3]{4\times 3}\]

OpenStudy (anonymous):

or \[10\sqrt[3]{12}\]

OpenStudy (saifoo.khan):

D for duck